In American football, the playing field is 53.33 yards (yd) wide by 120 yards (yd) long. For a special game, the field staff want to paint the playing field orange. Of course, they will use biodegradable paint available for purchase in 25-gallon (gal) containers. If the paint is applied in a thickness of 1.2 millimeters (mm) in a uniform layer, how many containers of paint will they need to purchase?


1
Expert's answer
2021-02-01T07:36:21-0500

"Solution:~playing ~field ~dimensions~

\\w~=~53.33~ 
\\yd= 53.33*0.9144=48.76~m

\\l=120~yd=109.728~m

\\container's ~ volume=25~gallon=0.094625 ~m^3

\\applied~ paints ~thickness ~(t)=1.2~mm=0.0012~m

\\so ~containers ~needed = \frac{playing~ field ~area ~X~ applied ~paint ~thickness}{volume ~of ~1 ~container}

 \\=\frac{48.76~X~109.728~X~0.0012}{0.094625}=\frac{6.42040}{0.094625} = 67.85 \approxeq 68 

\\\therefore Nearly ~68 ~containers ~needed."